OpenStudy (anonymous):

1.02^4t=3 I need to solve for t I so far have log4t(3)=1.02 I am not sure what to do next

hero (hero):

Take log of both sides to get 4t log(1.02) = log(3) Divide both sides by log(1.02) to get 4t = log(3)/log(1.02) Divide both sides by 4 to get t = log(3)/(4(log(1.02))

OpenStudy (anonymous):

how comes 4t moved out front?